Water drops fall at regular intervals from a roof at of 125 m from the ground. At an instant when 11th drop is about leave the roof and 1st drop is at a height of 45 m from the ground at that instant the dist of 10th drop from the roof is (g=10 m/s2)

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
DEAR STUDENT

Answer:

1.25m

Explanation:

as there are 11 drops there are 10 time intervals

total time =root  2h/g = root 2x250/g =5 sec

so each time interval=5/10=1/2sec

now the distance travelled by 10th drop  ; h of 10th drop=1/2gt2=1/2 x g x( 1/2)2 = 1.25m